OpenStudy (anonymous):

3% and 4% errors (Combinations of error) are given on a system, what is the total error?

OpenStudy (anonymous):

Total error is : 7x/100 I think so..

OpenStudy (anonymous):

I also think this, @waterineyes . Are there parameters which haven't been specified?

OpenStudy (anonymous):

Yes there are the parameters which have not been specified.. You are given errors percentage but you are not given how much the system is free from errors..

OpenStudy (kropot72):

The total error is given by taking the square root of the mean sum of the squares \[\sqrt{\frac{1.03^{2}+1.04^{2}}{2}}=\sqrt{1.07125}=1.035\] Total error is 3.5%

OpenStudy (anonymous):

error can be +ve or -ve.... sorry nothng left in the question, and this is not the answer too... this is the question from error analysis